OpenStudy (anonymous):

Which of the following is a solution to sqrt(x + 3) + 7 = 9

hero (hero):

sqrt(x + 3) + 7 = 9 sqrt(x + 3) = 2 x + 3 = 4 x = 4 - 3 x = 1

hero (hero):

Any questions?